# Think Less, Label Better: Multi-Stage Domain-Grounded Synthetic Data Generation for Fine-Tuning Large Language Models in Telecommunications

### v1

@@ I Introduction
We demonstrate the proposed framework in the context of telecom network troubleshooting, a domain where accurate solution plans are critical. In this setting, inputs may take the form of engineer-specified queries—for example, how to troubleshoot specific alarms or resolve a network faults. The framework outputs detailed, step-by-step troubleshooting solution plans that integrate evidence from multiple sources, including alarm logs, configuration management, performance counters, etc. The resulting synthetic QA dataset not only reflects the domain specificity and technical rigor of expert-authored responses but also significantly accelerates the preparation of high-qualith training data for RFT.
@@ III System Overview and Methods
The process begins with structured network fault alarms serving as prompts, along with topic documents that provide relevant background knowledge. A HippoRAG retriever module, powered by Qwen3-8B, identifies the top-3 relevant document chunks from the topic corpus using semantic similarity based on the built-knowledge graph. These retrieved chunks form the context for the initial generation of QA. The initial generation step uses a base model (Qwen-14B-Base) to synthesize raw QA pairs. This model is intentionally not instruct-tuned, allowing it to generate diverse and unbiased QA pairs. The outputs are then passed to an instruct-tuned model (Qwen3-8B/Qwen3-32B) with its ‘thinking’ mode disabled for efficiency. This model is provided with the top-3 supporting document chunks retrieved when HippoRAG is queried by the question, and rewrites the QA pairs for improved coherence, factual grounding, and clarity. The refinement emphasizes structured reasoning, which is essential for generating troubleshooting plans that follow logical step-by-step procedures. To ensure quality, we apply a filtering step using custom RAGAS scores, again powered by Qwen3-8B. Each QA pair is scored on key dimensions such as response groundedness, response relevancy, specificity, and aspect critic. Finally, only QA pairs above pre-defined thresholds are retained to form the RFT dataset.
@@ IV-A Seed Data Curation
We begin with seed data preparation to provide few-shot examples for both the base stage and refinement stage of domain-grounded synthetic data generation. We manually create 50 questions derived from troubleshooting documents covering diverse topics and alarms related to power systems in the telecom domain. For each question, we use HippoRAG in combination with LLMs (GPT-4o, LLaMA-3.3-70B-Instruct, and DeepSeek-R1-70B) to generate detailed stepwise solution plans. These QA pairs are then verified and refined by subject matter experts (SMEs) to ensure technical accuracy and completeness. Each finalized seed sample consists of three components: the question, the answer, and the top-3 relevant context chunks retrieved by HippoRAG.
@@ IV-B Domain-Grounded Synthetic Data Generation
Figure 2: Distribution of RAGAS metrics between Seed and Synthetic Dataset.
@@
Figure 4: Distribution of Pairwise Similarities between Seed and Synthetic Dataset.
Figure 5: Distribution of Pairwise Similarities between Refiner Qwen3-32B and Qwen3-8B in the Synthetic Dataset.
@@ IV-C Comparison of Base-Only, Instruct-tuned-only, and Hybrid Domain-Grounded Synthetic QA Generation
We compare the hybrid model (Qwen3-14B-Base + Qwen3-8B) against base-only (Qwen3-14B-Base) and instruct-tuned-only (Qwen3-8B) setups in multi-stage domain-grounded synthetic data generation. The hybrid approach not only produces more diverse questions but also delivers detailed, step-wise troubleshooting plans grounded in the top retrieved document chunks after RAGAS evaluation. As shown in Figures 6, and Table I, the base-only model exhibits higher density in the low-similarity region, which has avergae similarity score 0.35 compared with 0.65 in instruct-tuned-only, indicating greater question diversity. However, Table II reveals that most base-only outputs fail the RAGAS quality checks, leaving very few usable QA pairs for RFT. While, the hybrid model maintains an appropriate balance between diversity and generation ratio.
@@
Figure 6: Distribution of Pairwise Similarities with Base-only Model and Instruct-tuned-only Model in the Synthetic Dataset Generation.
@@ IV-E Runtime Performance
We optimize the runtime performance of our multi-stage domain-grounded synthetic data generation pipeline through several strategies. First, all Qwen3 models are served via vLLM, with LLM calls managed through LiteLLM and enhanced with retry logic, enabling robust multi-model support through configurable mappings. To ensure compatibility with evaluation, we implement a wrapper class that adapts vLLM’s inference engine to the BaseRagasLLM interface. For long-context reasoning, we configure RoPE scaling in vLLM using the YaRN strategy (factor = 2), allowing models to effectively process the top-3 retrieved context chunks with each question. With these optimizations, the full synthetic data generation process completes in approximately 45 minutes, followed by an additional 20 minutes for RAGAS evaluation to filter high-quality QA pairs suitable for RFT.
@@ V Conclusion
We introduced a multi-stage, domain-grounded pipeline for generating synthetic QA datasets to support RFT of LLMs. By combining retrieval via HippoRAG, diverse base generation, refinement for structured reasoning, and customized RAGAS evaluation, our approach produces diverse, high-quality, procedurally accurate troubleshooting data without human annotation. Experiments in telecom show that the hybrid pipeline outperforms base-only and instruct-tuned-only setups, while runtime optimizations enable efficient large-scale generation. This work demonstrates a scalable method to reduce reliance on expert labeling and highlights a practical path toward adapting LLMs for specialized, knowledge-intensive domains.

### v2

@@ I Introduction
We demonstrate the proposed framework in the context of telecom network troubleshooting, a domain where accurate solution plans are critical. In this setting, inputs may take the form of engineer-specified queries—for example, how to troubleshoot specific alarms or resolve a network faults. The framework outputs detailed, step-by-step troubleshooting solution plans that integrate evidence from multiple sources, including alarm logs, configuration management, performance counters, etc. The resulting synthetic QA dataset not only reflects the domain specificity and technical rigor of expert-authored responses but also significantly accelerates the preparation of high-quality training data for RFT.
@@ III System Overview and Methods
The process begins with structured network fault alarms serving as prompts, along with topic documents that provide relevant background knowledge. A HippoRAG retriever module, powered by Qwen3-8B, identifies the top-3 relevant document chunks from the topic corpus using semantic similarity based on the built-knowledge graph. These retrieved chunks form the context for the initial generation of QA. The initial generation step uses a base model (Qwen-14B-Base) to synthesize raw QA pairs. This model is intentionally not instruct-tuned, allowing it to generate diverse and unbiased QA pairs. The outputs are then passed to an instruct-tuned model (Qwen3-8B/Qwen3-32B) with its ‘thinking’ mode — i.e., explicit chain-of-thought decoding — disabled for efficiency. In this configuration, the model directly produces final answers without emitting intermediate reasoning tokens. This model is provided with the top-3 supporting document chunks retrieved when HippoRAG is queried by the question, and rewrites the QA pairs for improved coherence, factual grounding, and clarity. The refinement emphasizes structured reasoning, which is essential for generating troubleshooting plans that follow logical step-by-step procedures. To ensure quality, we apply a filtering step using custom RAGAS scores, again powered by Qwen3-8B. Each QA pair is scored on key dimensions such as response groundedness, response relevancy, specificity, and aspect critic. Finally, only QA pairs above pre-defined thresholds are retained to form the RFT dataset.
@@ IV-A Seed Data Curation
We begin with seed data preparation to provide few-shot examples for both the base stage and refinement stage of domain-grounded synthetic data generation. We manually create 50 questions derived from troubleshooting documents covering diverse topics and alarms related to power systems in the telecom domain. For each question, we use HippoRAG in combination with LLMs (GPT-4o, LLaMA-3.3-70B-Instruct, and DeepSeek-R1-70B) to generate detailed stepwise solution plans. These QA pairs are then verified and refined by subject matter experts (SMEs) to ensure technical accuracy and completeness. Each finalized seed sample consists of three components: the question, the answer, and the top-3 relevant context chunks retrieved by HippoRAG. Although the limited number of expert-curated seed QA pairs may constrain question diversity, they are used only as few-shot exemplars to steer the base model toward diverse QA generation, and retrieval-grounded RAGAS metrics with strict answerability and groundedness filtering ensure the robustness and quality of the resulting synthetic dataset.
@@ IV-B Domain-Grounded Synthetic Data Generation
Figure 2: Distribution of RAGAS metrics between Seed and Synthetic Dataset. Histograms are overlaid with transparency for direct comparison: blue represents seed dataset, red represents synthetic dataset, and overlapping regions appear as a blended color.
@@
Figure 4: Distribution of Pairwise Similarities between Seed and Synthetic Dataset. Histograms are overlaid with transparency for direct comparison: blue represents seed similarities, red represents synthetic similarities, and overlapping regions appear as a blended color.
Figure 5: Distribution of Pairwise Similarities between Refiner Qwen3-32B and Qwen3-8B in the Synthetic Dataset. Histograms are overlaid with transparency for direct comparison: blue represents Qwen3-32B, red represents Qwen3-8B, and overlapping regions appear as a blended color.
@@ IV-C Comparison of Base-Only, Instruct-tuned-only, and Hybrid Domain-Grounded Synthetic QA Generation
We compare the hybrid model (Qwen3-14B-Base + Qwen3-8B) against base-only (Qwen3-14B-Base) and instruct-tuned-only (Qwen3-8B) setups in multi-stage domain-grounded synthetic data generation. The hybrid approach not only produces more diverse questions but also delivers detailed, step-wise troubleshooting plans grounded in the top retrieved document chunks after RAGAS evaluation. As shown in Figures 6, and Table I, the base-only model exhibits higher density in the low-similarity region, which has average similarity score 0.35 compared with 0.65 in instruct-tuned-only, indicating greater question diversity. However, Table II reveals that most base-only outputs fail the RAGAS quality checks, leaving very few usable QA pairs for RFT. While, the hybrid model maintains an appropriate balance between diversity and generation ratio.
@@
Figure 6: Distribution of Pairwise Similarities with Base-only Model and Instruct-tuned-only Model in the Synthetic Dataset Generation. Histograms are overlaid with transparency for direct comparison: blue represents Base-only model, red represents Instruct-tuned-only model, and overlapping regions appear as a blended color.
@@ IV-E Runtime Performance
We optimize runtime performance through several system-level strategies. All Qwen3 models are served via vLLM, with inference managed by LiteLLM and augmented with retry logic to enable robust multi-model routing. For evaluation compatibility, we implement a wrapper that adapts vLLM’s inference engine to the BaseRagasLLM interface. To support long-context reasoning, we enable RoPE scaling in vLLM using the YaRN strategy (factor = 2), allowing efficient processing of the top-3 retrieved context chunks per query. With these optimizations, synthetic data generation completes in approximately 45 minutes, followed by an additional 20 minutes for RAGAS-based quality filtering.
@@ V Conclusion
We present a multi-stage, domain-grounded pipeline for synthetic QA generation to support RFT of LLMs. By integrating HippoRAG retrieval, diverse base generation, structured refinement, and customized RAGAS evaluation, the pipeline produces high-quality, procedurally grounded troubleshooting data without human annotation. Telecom experiments show that the hybrid approach outperforms base-only and instruct-tuned-only baselines, while runtime optimizations enable efficient large-scale generation. Overall, this work offers a scalable path to reducing expert labeling effort and adapting LLMs to knowledge-intensive domains.